### v1

# From User Preferences to Base Score Extraction Functions in Gradual Argumentation

Abstract.
Gradual argumentation is a field of symbolic AI which is attracting attention for its ability to support transparent and contestable AI systems. It is considered a useful tool in domains such as decision-making, recommendation, debate analysis, and others. The outcomes in such domains are usually dependent on the arguments’ base scores, which must be selected carefully. Often, this selection process requires user expertise and may not always be straightforward. On the other hand, organising the arguments by preference could simplify the task. In this work, we introduce Base Score Extraction Functions, which provide a mapping from users’ preferences over arguments to base scores. These functions can be applied to the arguments of a Bipolar Argumentation Framework (BAF), supplemented with preferences, to obtain a Quantitative Bipolar Argumentation Framework (QBAF), allowing the use of well-established computational tools in gradual argumentation. We outline the desirable properties of base score extraction functions, discuss some design choices, and provide an algorithm for base score extraction. Our method incorporates an approximation of non-linearities in human preferences to allow for better approximation of the real ones. Finally, we evaluate our approach both theoretically and experimentally in a robotics setting, and offer recommendations for selecting appropriate gradual semantics in practice.
Key words and phrases:
Gradual Argumentation; Base Score Extraction; User Preferences
1. Introduction
Abstract Argumentation Frameworks (AFs) Dung_JLP95 have long been known to be a promising tool for adaptable decision-making systems Amgoud_AAI09. These formalisms are composed of arguments and relations between them, representing conflicts. Semantics, whether extension-based or gradual, may then be used to determine the acceptance of the arguments, e.g., as potential decisions to be taken. However, in real-world scenarios involving human reasoning, having only arguments and their relations is rarely sufficient to align the decision towards human intentions. For example, some studies claim that the acceptability of arguments is related to the preferences of the audience to which they are addressed Kaci_CMA18.
Aligning decision-making systems with human preferences Gabriel_MM20, and also values Sierra_aamas19, has never been more important, given their increasing prevalence in humans’ daily lives. Within the field of Computational Argumentation, this has been accomplished by extending AFs into Preference-based AFs (PAFs) Amgoud_uai13, in which the arguments’ acceptability is based not only on their relations and the chosen semantics, but also on a preference ordering.
Figure 1. We introduce a methodology for converting a BAF, supplemented with a preference ordering of the arguments (top), into a QBAF (bottom). From the set of arguments and a user’s preference ordering, a Base Score Extraction Function $\nu$ is applied to the preference ordering to obtain the arguments’ base scores $\tau$ (middle), which allow for the QBAF to be used as normal for decision-making, i.e. comparing the decision arguments and selecting that with the highest strength. The output of the decision-making system is personalised to the user’s preferences.
[…]
A BAF appears at the top of the image, with a preference ordering over the arguments. In the middle of the image, a base Score Extraction Function is applied to the preference ordering to obtain the base scores of the arguments. At the bottom of the image, a QBAF is generated using the base scores that were obtained. A decision-making system selects an option between two topic arguments of the QBAF.
A current limitation in extension-based PAFs is that preferences are used to reduce or modify the framework Amgoud_IJAR14; Kaci_CMA18. Those reductions correspond to different intuitions, and each one is subject to criticism. In Battaglia_icsum24, reductions are avoided by extending PAFs to Quantitative Bipolar AFs (QBAFs) Baroni_IJAR19. In QBAFs, the arguments can attack and support each other, and each argument is assigned a base score, i.e. an intrinsic strength. Gradual semantics of QBAFs then assign a final strength to each argument, providing a different form of information than traditional AFs. QBAFs are currently used in several domains, such as decision support Battaglia_icsum24, product recommendation Rago_ijcai18, review aggregation Cocarascu_aamas19; Rago_AI25, decision-making reasoning Freedman_aaai25, given their potential for giving transparent and contestable systems Leofante_KR24; Yin_arxiv25. Nonetheless, in some scenarios, it is often unclear how to set the arguments’ base scores.
In this work, we propose a theoretically and experimentally motivated methodology for generating these base scores based on a given set of user preferences. Our approach is based on adding preferences to Bipolar Argumentation Frameworks (BAFs) Cayrol_ecsqaru05, i.e., AFs with an additional relation of support (but without a base score). We operate on the hypothesis that the given preference ordering allows for a base score extraction function to predict the arguments’ base scores. Although the focus of the paper is on converting BAFs into QBAFs, it could be applicable to AFs and support argumentation frameworks. Our approach is represented in Fig. 1. To ensure that our approach is able to represent realistic user preferences, on top of the usual I prefer $a$ over $b$, we introduce the possibility of having a preference relation in which one argument is much more preferred than another. We then provide some general axioms that any base score extraction function should satisfy. Next, we provide different design choices for adapting the base score extraction, which are formally characterised as potentially desirable, optional properties. Finally, we introduce some concrete base score extraction functions and evaluate them through both a theoretical analysis and experimentation on a proposed use case, assistive feeding in robotics.
Our contributions are as follows: (i) the introduction of base score extraction functions, for extracting the base scores of a set of arguments given a preference ordering, extended to non-linear preferences; (ii) the definition of a set of axioms and properties for base score extraction functions; (iii) the introduction of two concrete base score extraction functions; (iv) a theoretical and experimental evaluation of the two base score extraction functions.
To foster transparency and reproducibility, we have released the full implementation at https://github.com/acivit/From-Preferences-To-Base-Score-Extraction-Functions.
2. Preliminaries
In this section, we recall the definitions of BAFs, QBAFs and their gradual semantics.
A BAF Cayrol_ecsqaru05 is a triple $\mathcal{B}=\langle\mathcal{X},\mathcal{R}^{-},\mathcal{R}^{+}\rangle$, consisting of a finite set of arguments $\mathcal{X}$, a binary relation of attack $\mathcal{R}^{-}\subseteq\mathcal{X}\times\mathcal{X}$, and a binary relation of support $\mathcal{R}^{+}\subseteq\mathcal{X}\times\mathcal{X}$.
A QBAF Baroni_IJAR19 is a quadruple $\mathcal{Q}=\langle\mathcal{X},\mathcal{R}^{-},\mathcal{R}^{+},\tau\rangle$, where $\langle\mathcal{X},\mathcal{R}^{-},\mathcal{R}^{+}\rangle$ is a BAF and $\tau:\mathcal{X}\to[0,1]$ is a base score function. For any $a\in\mathcal{X}$, $\tau(a)$ is the base score of $a$.
In this work, we use the standard range for both the base scores and argument strengths, i.e., $[0,1]$, though others exist Amgoud_kr16. The strength of an argument $a\in\mathcal{X}$ is given by $\sigma(a)$, where $\sigma:\mathcal{X}\to[0,1]$ is a gradual semantics Baroni_IJAR19.
As mentioned, this work uses QBAFs for decision-making. In this context, the possible options of the decision process are included in the QBAF. These are called decision arguments. The QBAF is represented as a set of trees, with the root of each tree corresponding to a decision argument Rago_kr23. Let $\mathcal{Q}$ be a QBAF $\langle\mathcal{X},\mathcal{R}^{-},\mathcal{R}^{+},\tau\rangle$. For any arguments $a,b\in\mathcal{X}$, let a path from $a$ to $b$ be defined as a sequence of relations $(\gamma_{0},\gamma_{1}),$..., $(\gamma_{n-1},\gamma_{n})$ of length $n>0$, where $\gamma_{0}=a$ and $\gamma_{n}=b$, and, for any $1\leq i\leq n,(\gamma_{i-1},\gamma_{i})\in\mathcal{R}^{+}\cup\mathcal{R}^{-}$. Then, given a set of decision arguments $D\subseteq\mathcal{X}$, $\mathcal{Q}$ is a QBAF for $D$ iff i) $\nexists a\in\mathcal{X}\setminus\{D\}$ such that $\exists d\in D$ where $(d,a)\in\mathcal{R}^{+}\cup\mathcal{R}^{-}$ ii) $\forall a\in\mathcal{X}\setminus\{D\}$ there is a path from $a$ to at least one $d\in D$; and iii) $\nexists a\in\mathcal{X}$ with a path from $a$ to $a$. We leave cyclic frameworks to future work, as in Potyka_aaai21; Rago_kr23; Rago_AI25.
As a running example, we will use robot-assisted feeding. Here, a robot feeds a user who is unable to eat independently and must decide the pace of feeding, either slow or fast. The robot uses a QBAF drawing arguments from both the user and the robot’s observations, e.g., user status and history. The decision is made by selecting the decision argument with the higher strength. The arguments are: $(a)$ The patient gets bored if eating slowly, $(b)$ eating slowly does not stress the patient out, $(c)$ the patient is vulnerable and stress must be avoided, $(d)$ if the user does not eat fast, they will not have time with their niece who is coming to visit, $(e)$ the patient wants to tell their niece something important, and doing so will help the patient to feel more relaxed, and $(f)$ if the patient eats quickly, there is a (low) risk of dysphagia. The decision arguments are slow ($D_{1}$) and fast ($D_{2}$), which are possible feeding paces the robot can perform. The option selected is the one with the highest final strength. The framework is represented in Fig. 1.
3. Extracting Base Scores from Preference Orderings
In this section, we show that adding preferences to BAFs allows us to define a base score extraction function, which may exhibit some defined desirable axioms and properties. Then, the base score extraction function is used to extract the arguments’ base scores, effectively converting the BAF into a QBAF.
3.1. Adding Preferences to BAFs
Adding preferences over arguments is a common approach to personalising argumentation frameworks for users Modgil_AI13; Amgoud_uai13; Mailly_kr20. We introduce the formal definition of preferences over arguments:
Definition 1.
Given a set of arguments $\mathcal{X}$, a preference $\succeq_{p}$ over $X$ is a reflexive and transitive relation on $\mathcal{X}$. Given $\in\mathcal{X}, aba\succeq b$ and $a\npreceq b$ will be denoted as $a\succ b$, while $a\preceq b$ and $a\succeq b$ will be denoted as $a\simeq b$.
Setting the base scores according to the preference ordering of a BAF’s arguments in a decision-making environment can give an intuition of the most suitable decision in line with a user’s preferences. The following examples show the importance of properly setting the arguments’ base scores:
Example 0.
Consider the framework from Fig. 1. If the base scores are arbitrarily set to 0.5, giving equal importance to all arguments, using the Quadratic Energy Potyka_kr18 model, the final strengths of the decision arguments are $\sigma(D_{1})=0.5$ for slow, and $\sigma(D_{2})=0.505$ for fast, then the robot would choose to move fast.
Now imagine a scenario where a user gives more importance to the arguments related to its safety ($c$ and $f$), less importance to the arguments related to its relaxation ($b$ and $e$), and the least importance to the arguments related to its enjoyment ($a$ and $d$):
Example 0.
(cont) The resulting preference ordering is: $c\simeq f\succ b\simeq e\succ a\simeq d$. Intuitively, the base scores of $c$ and $f$ should be greater than those of $b$ and $e$, and these greater than the base scores of $a$ and $d$. For example, they could be set to $\tau(c)=\tau(f)=0.75$, $\tau(b)=\tau(e)=0.5$, and $\tau(a)=\tau(d)=0.25$. Here, using the Quadratic energy model, the final strengths are $\sigma(D_{1})=0.54$ and $\sigma(D_{2})=0.45$, then the robot would choose to move slowly, which is the opposite of what has been selected with equal base scores.
3.2. Adapting to Non-Linear Human Preferences
Human preference relations are generally not linear Northcraft_OBHDP98; people’s choices and priorities often fluctuate instead of increasing or decreasing in a simple, linear manner as conditions change. This non-linearity makes it complex to set the base scores properly. A possible approach to including these non-linearities could be achieved by allowing the user to order their preferences gradually within a range. However, setting a gradual score is usually slower and has more sample noise than cardinal ordering Yannakakis_acii11; Shah_arxiv14.
Another approach, which is used in this work, is to approximate the non-linearities by introducing a new preference relation, where a user may have a much greater ($\succ\!\succ$) preference for one argument over another. This novel relation considers the non-linearity and is easy to recognise from human feedback. Definition 1 is extended as follows:
Definition 2.
Given $a,b\in\mathcal{X}$, if $a$ is much more preferred than $b$, the preference relation will be denoted as $a\succ\!\succ b$. This relation is transitive and satisfies $a\succ b$. Therefore, a preference relation between arguments $a,b\in\mathcal{X}$ in a preference ordering $\succeq_{p}$ can be:
  • $a\succ b$, which denotes strict preference;
  • $a\simeq b$, which denotes indifference;
  • $a\succ\!\succ b$, which denotes a much stronger preference.
Example 0.
Given the arguments $a,b,c\in\mathcal{X}$, with a preference ordering: $a\succ\!\succ b\succ c$. Here, we may expect that $\tau(a)$ is much greater than $\tau(b)$ and $\tau(c)$, e.g., 0.9, 0.3, and 0.1, respectively.
This new relation breaks the linearity in the preference ordering, providing a more realistic approximation to human preferences, since it has a greater expressiveness.
3.3. Base Score Extraction Functions
Although PAFs and BAFs offer means to capture the evaluation of arguments, they remain categorical and context-sensitive. In contrast, integrating a method to extract base scores from preference orderings would provide a more granular and objective evaluation mechanism. Furthermore, the literature on QBAFs offers methods to provide transparent and counterfactual explanations Kampik_IJAR24; Yin_kr24 and allows for fine-tuning the framework to better adapt to users, which is crucial in some decision-making systems, such as robotics.
Therefore, we introduce base score extraction functions that compute the arguments’ base scores given a preference ordering.
Definition 3.
A Base Score Extraction Function $\nu$ is a function that takes a set of arguments $\mathcal{X}$ and a set of preference orderings $\mathcal{P}$ to obtain a set of all the base score functions $\mathcal{T}$, therefore $\nu:\mathcal{X}\times\mathcal{P}\rightarrow\mathcal{T}$.
Obtaining the base scores of a set of arguments allows the transformation of BAFs into QBAFs. To simplify the formalisations in the following sections, we consider $\nu_{P}(\mathcal{X})$ as the base score extraction function $\nu$ applied to $\mathcal{X}$ with a given preference ordering $\succeq_{p}$.
3.4. Desirable Properties of Base Score Extraction Functions
In the following, we present the different axioms and properties of the base score extraction functions. First, we define the Preference Coherence axiom, which states that if one argument is preferred over another, the former’s base score will be greater.
Axiom 1.
(Preference Coherence) Given a set of arguments $\mathcal{X}$ with arguments $a,b\in\mathcal{X}$ and a preference ordering $\succeq_{p}$, a base score extraction function $\nu_{p}(\mathcal{X})$ satisfies Preference Coherence iff:
  • if $a\succ b$ or $a\succ\!\succ b$ then $\tau(a)>\tau(b)$; and
  • if $a\simeq b$ then $\tau(a)=\tau(b)$.
We now define Preference Relation Coherence, which provides a differentiation between the preference relations presented. Intuitively, if an argument is much preferred ($\succ\!\succ$) over another, the difference of their base scores will be greater than if the preference relation is only preferred ($\succ$).
Axiom 2.
(Preference Relation Coherence) Given a set of arguments arguments $a,b,c,d\in\mathcal{X}$ and the preference ordering $a\succ\!\succ b\succ c\simeq d$, a base score extraction function $\nu$ satisfies Preference Relation Coherence iff $\tau(a)-\tau(b)>\tau(b)-\tau(c)>\tau(c)-\tau(d)$.
In our approach, the focus on extracting the base scores is set on the preference structure rather than the comparison between elements. The following axiom is necessary to define whether two or more preference orderings are isomorphic. If the preferences of two sets of arguments exhibit the same relationships, they can be said to have isomorphic preferences.
Definition 4.
Let $\succeq_{p}$ and $\succeq_{p^{\prime}}$ be two preference orderings over two identical sets of arguments, $\mathcal{X}$ and $\mathcal{X}^{\prime}$, respectively. Those preference orderings are isomorphic iff there is a bijective function $f:\mathcal{X}\to\mathcal{X^{\prime}}$ such that $\forall i,j\in\mathcal{X},\ i\succeq_{p}j\Leftrightarrow f(i)\succeq_{p^{\prime}}f(j)$ [^1] We omit the subscript of the preference relation when unambiguous.. It is denoted as $\succeq_{p}\simeq\succeq_{p^{\prime}}$.
We require that a base score extraction function preserves the structural distinctions of users’ preferences. If two users have different (non-isomorphic) preference orderings, the assigned base scores must reflect these differences.
Axiom 3.
(Preference Structure Coherence) Given two preference orderings that are not isomorphic $\succeq_{p}\not\simeq\succeq_{p^{\prime}}$, a base score extraction function $\nu$ satisfies Preference Structure Coherence iff for a set of arguments $\mathcal{X}$, then $\nu_{p}(\mathcal{X})\neq\nu_{p^{\prime}}(\mathcal{X})$.
Note that if two sets of arguments have different base scores, they will not necessarily have different preference ordering.
3.5. Design Choices for Base Score Extraction Functions
The presented axioms offer the necessary conditions that the base score extraction functions must satisfy. Despite that, determining the base scores is still a flexible and context-dependent challenge. This section determines different design choices that must be considered when designing a base score extraction function.
3.5.1. Setting the Range
A choice that must be made is considering where the base scores of the most and least preferred arguments should be placed. Establishing them in 1 and 0, respectively, determines a strict and high importance of the preferences, but also allows for possible saturations from incoming supports or attacks, which would provoke information loss, an existing concern in PAFs. For a further investigation, we define the base score limits.
Definition 5.
[…]
Definition 6.
[…]
The range is constrained within the range of the gradual semantics, i.e. $\mathbb{I}_{\sigma}\in[0,1]$. It must be satisfied that $0\leq\bot\leq\top\leq 1$. A base score extraction function will be considered normalised if its limits are set to $\top=1$ and $\bot=0$.
Property 1.
(Base Score Normalisation) A base score extraction function $\nu$ satisfies Base Score Normalisation iff $\top=1$ and $\bot=0$.
In some scenarios, the limit base scores might be centred at 0.5 to differentiate between important and less important arguments.
Property 2.
(Base Score Centralisation) A base score extraction function $\nu$ satisfies Base Score Centralisation iff $\top=1$-$\bot$.
The range influences the base scores compression. For example, if the range is too low, the base scores would be very similar. Hence, the preferences of the arguments would not be properly reflected. The following example represents the concern.
Example 0.
Given the arguments $a,b,c,d,e,f\in\mathcal{X}$, and the preference ordering: $c\simeq f\succ b\simeq e\succ a\simeq d$. If the limits were set to $\top=0.9$ and $\bot=0.85$, the base scores could be placed only between $[0.85,0.9]$ (e.g., $\tau(c)=\tau(f)=0.9,\tau(b)=\tau(e)=0.875,\tau(a)=\tau(d)=0.85$). Alternatively, if $\bot=0.1$, the base scores could be placed in the range $[0.1,0.9]$ (e.g., $\tau(c)=\tau(f)=0.9,\tau(b)=\tau(e)=0.5,\tau(a)=\tau(d)=0.1$).
Although the base scores from the previous examples satisfy all the axioms, in the first case, they may not differentiate the preferences enough for some applications, since the difference between them is too low. As seen in the second case, decreasing the bottom limit results in a more expanded range of base scores in which the preference relations are more relatable.
3.5.2. Preference Regularity
In many real-world cases, users express their preferences in a consistent or regular way, meaning that the perceived difference in importance between two consecutively preferred arguments is roughly uniform across the ordering. For instance, if a user slightly prefers argument $a$ over $b$, and $b$ over $c$, the gap in importance between $a$ and $b$ is comparable to that between $b$ and $c$. Assuming such regularity simplifies the computation of base scores.
Before introducing the property, it is necessary to define the concept of adjacent arguments.
Definition 7.
Let $a,b\in\mathcal{X}$ be two arguments, and $\succeq_{p}$ a preference ordering over the arguments in $\mathcal{X}$, the arguments $a,b$ are adjacent in the preference ordering iff one of the following conditions holds:
  • $a\succ b$ and $\nexists c$ such that $a\succ c\succ b$;
  • $a\succ\!\succ b$ and $\nexists c$ such that $a\succ\!\succ c\succ b$ or $a\succ\!\succ c\succ\!\succ b$ or $a\succ c\succ\!\succ b$;
  • $a\simeq b$.
Property 3.
(Base Score Relation Regularity) A base score extraction function $\nu$ satisfies Base Score Relation Regularity iff for a given set of arguments $\mathcal{X}$, for any preference relation $p\in\{\succ\!\succ,\succ,\simeq\}$, and any two pairs of adjacent arguments in the preference ordering, $(a,b)$ and $(c,d)$, with the same preference relation, then $\nu_{p}(a)-\nu_{p}(b)=\nu_{p}(c)-\nu_{p}(d)$.
Example 0.
(cont.) Given the previous arguments $a,b,c,d,e,f\in\mathcal{X}$, and the preference ordering: $c\simeq f\succ b\simeq e\succ a\simeq d$. If the limits were $\top=0.9$ and $\bot=0.1$, if the base score extraction function satisfies Base Score Relation Regularity, then the base scores would be $\tau(c)=\tau(f)=0.9,\tau(b)=\tau(e)=0.5,\tau(a)=\tau(d)=0.1$. Otherwise, the base scores $\tau(b)$ and $\tau(c)$ could be arbitrarily placed between 0.1 and 0.9.
If this regularity condition is not enforced, one could theoretically encode preference intensity through irregular spacing between arguments rather than using a separate much greater relation. Yet, this flexibility makes the base score extraction process more difficult, as there is no clear rule for assigning consistent numerical differences.
Another property relating to the regularity of the structure is that if new arguments are added to the framework, and these are equally preferred to existing ones, the base scores of the rest of the arguments will not change.
Property 4.
(Base Score Preference Stability) Let $\succeq_{p}$ be a preference ordering for arguments in $\mathcal{X}$, and a new set of arguments $S$ such that $\mathcal{X}^{\prime}=\mathcal{X}\cup S$, a base score extraction function $\nu$ satisfies Base Score Preference Stability iff for the new preference ordering $\succeq_{p^{\prime}}$ such that $\forall s\in S$, $\exists a\in\mathcal{X}$ such that $a\simeq s$, then $\nu_{p}(\mathcal{X})=\nu_{p^{\prime}}(\mathcal{X}^{\prime}\setminus S)$.
3.5.3. Preferences ratios
While Preference Regularity assumes that all consecutive preferences correspond to equal differences in base scores, this may not always reflect how people actually weigh options. In practice, users often express varying intensities in their preferences. For instance, one argument may be only slightly preferred over another, while another is much more important. To model such heterogeneous gaps, we introduce the concept of preference ratios, which quantify how much stronger a much greater preference should be compared to a regular preference ($\succ$). This ratio provides a systematic way to translate qualitative intensity (e.g., “much more important”) into quantitative differences between base scores.
Example 0.
(cont. ) Now the preference ordering becomes: $c\simeq f\succ\!\succ b\simeq e\succ a\simeq d$. From asking the human, the ratio of the much greater preference is set to 3. The base score extraction function limits are set to 0.9 and 0.1. Then, the base scores could be set to $\tau(c)=\tau(f)=0.9,\tau(b)=\tau(e)=0.3,\tau(a)=\tau(d)=0.1$. Here, the difference between preferred arguments is 0.2, while for much preferred arguments is 0.6, hence the ratio is respected.
Observe that the ratio between the different preferred relations and the limits must be placed carefully, as there is a risk of undermining the preferred or much more preferred relations. The following example shows two scenarios where this might occur:
Example 0.
Consider $a,b,c,d\in\mathcal{X}$, with the preference ordering: $a\succ b\succ\!\succ c\succ d$. If the ratio is set to 1.33, with the limits at 0.75 and 0.25, the base scores would be $\tau(a)=0.75,\tau(b)=0.6,\tau(c)=0.4,\tau(d)=0.25$. The difference between preferred arguments is 0.15, while it is 0.2 for much more preferred arguments. At first sight, it does not seem representative enough. Alternatively, setting a ratio of 96 and the limits at 0.99 and 0.01, the base scores would be $\tau(a)=0.99,\tau(b)=0.98,\tau(c)=0.02,\tau(d)=0.01$. Here, the difference between the adjacent preferred arguments is 0.01, which intuitively may not be representative in many real-world contexts.
These examples highlight that the preference ratio serves as a tuning parameter controlling how sensitively the system reacts to much greater preferences. A well-chosen ratio helps capture meaningful differences in user priorities while maintaining numerical stability in base score extraction. The combination of these three different design choices thus gives a formal representation of how base score extraction functions may be personalised to different settings.
4. Concrete Base Score Extraction Functions
In this section, two base score extraction functions are introduced. Those functions are based on the preference ordering of the arguments and allow for the presented design choices. An algorithm is developed to compute the base scores. Some examples of its use are shown using the running example. By definition, as the base score extraction functions are based on preference orderings, then, scenarios with only one argument are not considered.
Since preference orderings are monotonic, considering a descending preference ordering (from the most preferred to the least preferred), the base score extraction function should also be monotonically decreasing.
Definition 8.
A base score extraction function is monotonically decreasing (increasing) iff given two arguments $a,b$ where $a\succeq b$, then $\nu_{p}(a)\leq\nu(b)$ ($\nu_{p}(a)\geq\nu(b)$, respectively).
Proposition 0.
A monotonically increasing base score extraction function violates Axiom 1 for a descending preference ordering. [^4] The proof of this proposition is found in Appendix A.
Creating a monotonically decreasing function directly from the preference ordering is possible. The following method consists of assigning a distance $d$ to each argument according to the preference ordering and then normalising it between the desired range. The distance is relative to the most preferred argument.
Base Score Extraction Function 1.
(Adaptable range) Our first base score extraction function allows for the range to be set and is defined as follows:
$\nu_{1}(\top,\bot,D,d)=\bot+(\top-\bot)\cdot\frac{D-d}{D-1}$ (1)
where $D$ is the highest distance, $d$ is the current distance, $\bot$ is the base score of the least preferred argument, and $\top$ is the base score for the most preferred argument.
The edges of this base score function correspond to when $d=1$ and $d=D$, in which the function becomes $v_{1}=\top$ and $v_{1}=\bot$ respectively. Then, with the restriction $0\leq\bot\leq\top\leq 1$, both values are ranged in $[0,1]$. Given these constraints, the function is defined inside $[0,1]$.
The next base score extraction function consists of setting the compression of the base scores and the distances towards the edges.
Base Score Extraction Function 2.
(Adaptable squeezing and distancing) Our next base score extraction function allows a parameter to determine how compressed the base scores are to be set, and another to determine how distant the edges are, and is defined as:
$\nu_{2}(\alpha,\beta,D,d)=\frac{D-d+\alpha}{D-1+\beta}$ (2)
where $\beta\in[0,+\infty)$ is a constant related to the compression of the base scores, and $\alpha\in[0,+\infty)$ is a constant related to the distance to the edges. Intuitively, the bigger $\beta$ is, the more compressed the base scores are, while the bigger $\alpha$ is, the closer the preferred argument will be to 1.
It is necessary that $\alpha\leq\beta$, otherwise the base scores would be greater than 1. With $\beta=D-1$, the spread of the base scores is compressed in a range of 0.5. This base score extraction function is more suitable than the first one when the number of arguments is unknown or very likely to change. The first one allows for the range to be set directly, which provides more control over the edges.
Algorithm 1 shows how to extract the base scores from a set of arguments $\mathcal{X}$ and a preference ordering $\succeq_{p}$. First, the algorithm assigns $d=1$ to the most preferred argument, then, for the following arguments in the ordering, $d$ increases by $\delta$ between preferred relations and by $\Delta$ for the much preferred relationship.
Definition 9.
The distance between preferred arguments is $\delta\in(0,\infty)$, and for much more preferred arguments is $\Delta\in(0,\infty)$.
After assigning this distance, the base scores are obtained using one of the proposed base score extraction functions. An example applying these distances in Algorithm 1 is shown:
Algorithm 1 Flexible Normalised Base Score Extraction
1:A set of arguments $\mathcal{X}$, its preference ordering $\succeq_{p}$, the increase for greater preference relation $\delta$, for much greater $\Delta$, the constants $c_{1},c_{2}$ ($\top,\bot$ or $\alpha,\beta$) depending on the base score extraction function $\nu(c_{1},c_{2},D,d)$.
2:Set of argument base scores $\mathcal{T}$
[…]
8:for each argument $a\in\mathcal{X}$ ordered by $\succeq_{p}$ do
9:  if $a$ is not the most preferred then
10:   $b\leftarrow$ the previous argument in $\succeq_{p}$
11:   if $b\succ a$ then
12:     $d\leftarrow d+\delta$
13:   else if $b\succ\!\succ a$ then
14:     $d\leftarrow d+\Delta$
15:   else
16:     $d\leftarrow d$
17:  $\tau(a)\leftarrow\nu(c_{1},c_{2},D,d)$
18:  $\mathcal{T}\leftarrow\tau(a)$
19:return $\mathcal{T}$
Example 0.
[…]
The base scores obtained can be used for decision-making, as shown in the following example.
Example 0.
(cont.) The final strengths for the decision arguments are computed based on the previous example base scores, using the Quadratic Energy Model Potyka_kr18, which results in: $\sigma(D_{1})=0.54$ and $\sigma(D_{2})=0.40$, then the robot decides to move slowly. Note that the user prefers arguments $c$ and $f$, which increase the strength of the slow option and decrease the strength of the fast option, respectively. The arguments $b$ and $e$ increase and decrease the strength of the slow option, respectively. Finally, the arguments $a$ and $d$ aim to make the robot go fast. The algorithm’s selection of moving the robot slowly accurately adapts to the user’s preferences.
These base score extraction functions offer different benefits. First, both are flexible, allowing the base scores to be set in a desired range and proportionally adjusting the difference of base scores between arguments. Additionally, the parameters are easily modifiable, allowing corrections and learning from human feedback. Other functions could be used, e.g., an exponentially increasing or decreasing function, but there is a high chance that they do not adjust properly to a user’s preferences. For example, given five arguments $a,b,c,d,e\in\mathcal{X}$ ordered as: $a\succ b\succ c\succ d\succ e$, an exponentially decreasing function could set the base scores to $\tau(a)=1$, $\tau(b)=0.9$, $\tau(c)=0.75$, $\tau(d)=0.5$, and $\tau(e)=0.1$. Assuming that the distribution of the base scores follows that function is most probably incorrect. Allowing the user to confirm the ordering as: $a\succ b\succ c\succ d\succ\!\succ e$, the base scores could be set to $\tau(a)=1$, $\tau(b)=0.866$, $\tau(c)=0.677$, $\tau(d)=0.5$, $\tau(e)=0.1$. These base scores are similar to the previous ones, but the confirmation of the large gap between arguments $d$ and $e$ provides a more accurate representation of the real preferences.
5. Evaluation
This section presents a theoretical validation of the concrete base score extraction functions based on how they satisfy the theoretical properties we defined. We undertake an experimental evaluation, showing the pros and cons of each base score extraction function. Finally, we give an intuition of which gradual semantics to select depending on the decision-making context.
5.1. Theoretical Analysis
First, we show under which conditions our proposed base score extraction functions satisfy the axioms and properties presented. This is summarised in Table 1. The proofs can be found in Appendix A.
$\nu$ A1 A2 A3 P1 P2 P3 P4
[…]
Table 1. Summary of the conditions for satisfying the presented axioms (A) and desirable properties (P).
For Axiom 1, constants $\Delta$ and $\delta$ must be greater than zero.
Proposition 0.
Setting the constants $\Delta>0$ and $\delta>0$ satisfies Axiom 1 for both base score extraction functions.
Preference ordering $\succeq_{p}$ Design Choices QE Strengths EB Strengths DF Strengths
$\top$ $\bot$ $\Delta/\delta$ $\sigma_{slow}$ $\sigma_{fast}$ $\sigma_{slow}$ $\sigma_{fast}$ $\sigma_{slow}$ $\sigma_{fast}$
$a\simeq b\simeq c\simeq d\simeq e\simeq f$ - - - 0.5 0.51 0.50 0.52 0.44 0.63
$c\simeq f\succ\!\succ b\simeq e\succ a\simeq d$ 0.9 0.1 3 0.58 0.33 0.56 0.39 0.78 0.23
0.9 0.1 5 0.57 0.32 0.55 0.39 0.79 0.2
0.75 0.25 5 0.52 0.4 0.53 0.43 0.63 0.38
$b\simeq e\succ a\simeq d\succ\!\succ c\simeq f$ 0.8 0.2 3 0.5 0.63 0.52 0.60 0.28 0.87
0.6 0.4 3 0.5 0.53 0.5 0.54 0.40 0.71
$a\simeq d\succ\!\succ c\simeq f\succ b\simeq e$ 0.8 0.2 4 0.36 0.6 0.4 0.59 0.15 0.76
1 0 4 0.25 0.7 0.37 0.65 0 0.9
Table 2. Examples of preference orderings with different design choices for the base score extraction function 1. The value in bold between $\sigma_{slow}$ and $\sigma_{fast}$ is the system output. The (-) symbol indicates when a design choice is not needed, which is the case where all arguments are equally preferred, and their base scores are set to 0.5.
For Axiom 2, it is necessary that the distance between much more preferred arguments is bigger than the distance between preferred arguments, and must be greater than zero.
Proposition 0.
Setting $\Delta>\delta>0$ satisfies Axiom 2 for both base score extraction functions.
Since the base score extraction functions are based on the distance between arguments given an ordering, if that ordering is changed, the base scores’ ordering will also change.
Proposition 0.
Both base score extraction functions satisfy Axiom 3.
[…]
Proposition 0.
[…]
Proposition 0.
[…]
Proposition 0.
[…]
Proposition 0.
[…]
To have regular preference relations, the variables of the algorithm to compute the distance ($\Delta,\delta$) must be constant.
Proposition 0.
[…]
Property 4 determines that the base scores of a set of arguments with a given preference relation remain constant when adding a new argument that is equally preferred to an existing one.
Proposition 0.
Both base score extraction functions satisfy Base Score Preference Stability.
Finally, if the selected gradual semantics is monotonic and balanced Baroni_IJAR19, if an argument is preferred to another and both have the same influences, the preferred argument will be stronger.
Proposition 0.
Given a base score extraction function $\nu$, where $\nu$ satisfies Axiom 1, and a semantics $\sigma$, where $\sigma$ satisfies monotonicity and balance, for any arguments $a$, $b$ where $a\succ b$ and $a$ and $b$ have equivalent supporters and attackers, then $\sigma(a)>\sigma(b)$.
5.2. Experimentation
The aim of this experiment is to evaluate the practical behaviour of the proposed BSEFs under different user preference profiles and design choices, and to examine how various gradual semantics influence the resulting decisions. Specifically, we test whether the generated QBAFs lead to consistent, interpretable, and preference-aligned outcomes across semantics.
We consider the running example from Fig. 1 in which the robot must choose between slow ($D_{1}$) and fast ($D_{2}$) feeding paces (see Example 3.1). Each QBAF instance is constructed using six arguments and a pair of decision arguments.
To explore the robustness of our approach, we generated 30,000 random samples of preference orderings and design choices, recording the option selected by three gradual semantics models. These models are: the Quadratic Energy (QE) Potyka_kr18, the Euler-Based (EB) Amgoud_IJAR18, and the DF-QuAD (DF) Rago_kr16, which are among the most used in the literature. The decision arguments’ base scores are set to 0.5. We computed pairwise agreement and Cohen’s Kappa coefficient ($\kappa$) to assess consistency between models. Under the centralisation property, the pairwise agreement values were QEEB=0.98QE-EB=0.98, QEDF=0.85QE-DF=0.85, and EBDF=0.84EB-DF=0.84. The corresponding Cohen’s Kappa scores were QEEB=0.96QE-EB=0.96, QEDF=0.65QE-DF=0.65, and EBDF=0.62EB-DF=0.62. Without centralisation, the pairwise agreement values were QEEB=0.96QE-EB=0.96, QEDF=0.81QE-DF=0.81, and EBDF=0.81EB-DF=0.81, with Cohen’s Kappa scores of QEEB=0.92QE-EB=0.92, QEDF=0.52QE-DF=0.52, and EBDF=0.52EB-DF=0.52. These results show a strong alignment between QE and EB, while DF-QuAD diverges more, particularly when base score ranges are extreme (e.g., $\tau(e)\approx 1$, $\tau(e)\approx 0$). This difference shows the structural properties of the methods discussed in Sec. 5.3.
To further analyse the influence of user preferences and design parameters, Table 2 reports examples of preference orderings combined with different design choices, evaluated under the three gradual semantics. The preference relations are considered regular to avoid complexity in experimentation and noise in the results.
The results confirm that the preference ordering itself is the dominant determinant of the final decision, while design choices act as fine-tuning factors that adjust numerical sensitivity but sometimes invert the selected option. For instance, when arguments related to safety ($c$ and $f$) are prioritised over those concerning enjoyment ($a$ and $d$), all semantics consistently select the slow feeding pace, aligning with the user’s intended caution. Conversely, when enjoyment-related arguments are most preferred, the system reliably opts for fast feeding, demonstrating the coherence of the extracted base scores with user values.
Design choices such as the range limits affect the magnitude of the option strengths rather than the decision direction. Narrower ranges (e.g., $(\top,\bot)=(0.75,0.25)$) compress the difference between argument strengths, producing less decisive but more stable outcomes, while broader ranges (e.g., $(1,0)$) amplify contrasts and make the system more sensitive to preference changes. Similarly, the preference ratio $(\Delta/\delta)$ controls how much stronger a much greater preference influences the base scores. Increasing this ratio enhances differentiation between preference tiers but may also overemphasise extreme preferences, especially under DF-QuAD, which is more responsive to high base score values.
5.3. Discussion on the Gradual Semantics
The previous analysis revealed that, although all semantics preserve preference–decision coherence, their sensitivity to base score variations differs substantially. This behaviour stems from the aggregation and influence functions each model employs.
Figure 2. Influence of an argument’s base score and its supporter or attacker on the final argument strength under different gradual semantics. Each subplot corresponds to a distinct semantics: DF-QuAD (left), Quadratic Energy Model (centre), and the Euler-based Semantics (right). The x-axis represents the base score of the influenced argument, and the y-axis its resulting final strength. Blue lines denote when the influencing argument is a supporter (+), while red lines denote when it is an attacker (-). Each line has assigned its aggregation ($\alpha$), with colour intensity indicating its strength.
[…]
Each subplot shows the influence of an argument’s base score (x-axis), and the different aggregation strengths (lines) for that argument’s final strength (y-axis) for the different gradual semantic models. The left figure is for DF-QuAD, the centre one is for Quadratic Energy, and the right is for Euler-based. The blue lines are for supports, while the red lines are for attacks. At the tip of the lines is written the base score of the supporter (+) or attacker (-). The colour intensity of the lines represents the strength of the attack or support.
Figure 2 shows the influence of a single attacker or supporter (the influencer), and an argument’s base score (the influenced) for its final strength. The DF-QuAD semantics use product aggregation with a linear influence function, which causes the influencer’s base score to dominate the final outcome. When an attacking (or supporting) argument has a base score close to 1, it almost completely drives the influenced argument’s strength toward 0 (or 1), regardless of the influenced base score. This strong reactivity can be desirable in safety-critical scenarios, where highly preferred arguments should decisively determine the outcome, but it may reduce robustness in settings with uncertain or noisy preferences. In contrast, the QE combines sum aggregation with a 2-Max influence function. Here, both the influencer and the influenced arguments contribute comparably to the final strength, leading to smoother transitions and more gradual adaptation to preference changes (e.g., given an influencer with a base score of 1, the influenced’s final strength will be $\sigma=0.5\tau+0.5$ if it receives a support, and $\sigma=0.5\tau$ if it receives an attack). The QE, therefore, offers a good compromise between responsiveness and stability, which makes it particularly suitable for explainable decision-support systems that require consistent reasoning traces. Finally, the EB is the most conservative, using sum aggregation with an Euler-based influence function. When an argument’s base score is close to the extremes (0 or 1), its final strength remains nearly fixed at those values, regardless of new supports or attacks. This behaviour preserves high-confidence arguments, making EB appropriate when users insist on arguments being non-negotiable or when decisions must respect rigid value priorities.
6. Related Work
Gradual argumentation is highly reliant on the arguments’ base scores, and its literature has unveiled different ways of getting those, depending on the context. First, in Baroni_AC15, the authors presented the QuAD framework, intending to support debates on design alternatives. Since those scenarios do not involve a large number of voters to compute a base score, the recommended approach is to have an expert assess the base score, or obtain it from a given set of criteria and their importance. As an extension, in Rago_icppmas17, they presented QuAD-V, which allows users to vote in favour of or against arguments. The votes are used to compute a base score for each argument according to multiple users. A similar procedure is done in Tarle_aamas22. In another work Cocarascu_aamas19 in which a QBAF is generated from movie reviews to extract the most relevant information, the base scores are obtained from an aggregation of critics’ votes. Meanwhile, within structured argumentation, the work in Tamani_icfs14 is focused on ASPIC argumentation frameworks with fuzzy set theory and uses the expertise knowledge to adjudicate the importance of an argument in an environment, similarly to a base score. Another use of QBAFs is to apply their methods for explainable AI, mainly transforming (deep) neural networks into QBAFs. For instance, in Albini_arxiv20, the base scores of the arguments are obtained from the attackers and supporters and activations in the network, similarly to the work in Potyka_aaai21. Differently from the previous approaches, the work from Battaglia_icsum24 sets the base scores based on users’ preferences, similarly to this work. The preferences are integrated to personalise decision support. In a nutshell, the initial value of the base scores is arbitrarily set, and then a discount factor is applied to the less preferred arguments. This work proposes a method to use preferences to determine the initial value of the base scores.
In summary, two main limitations are found in existing work. First, most approaches do not consider the preferences of a single user to set the base scores, making it difficult to personalise argumentation-based decision-making systems. Second, when preferences are considered, the base scores are arbitrarily initialised and then adapted based on the preferences. In this work, we address those research gaps.
7. Conclusions and Future Work
This work introduces base score extraction functions, a principled method to derive argument base scores from user use preference orderings in gradual argumentation. We defined a set of axioms and desirable properties, proposed two concrete base score extraction functions with tunable design choices, and incorporated non-linear preference intensity to better capture human reasoning. Theoretical analysis confirmed that the functions satisfy coherence and regularity properties, while experiments in an assistive robotics scenario showed that preference ordering dominated the decision outcome and design choices primarily modulate sensitivity.
Future directions include the fine-tuning or definition of the different design choices given the context and user feedback, allowing partial orderings in preferences in case some arguments cannot be compared Visser_wtafa11, and computing the aggregation when using this approach when considering more than one user’s preferences in the decision-making Rago_AI25. A philosophical question which should be commented on is whether it is realistic to consider the scenario in which the most preferred argument supports the least preferred argument, which is not considered in this work. Finally, the extension of this work into value-based argumentation could provide an approach for more value-aligned decision-making Atkinson_AL21; Bodanza_AC23. This could be potentially achieved by grouping the arguments into sets of values and ordering the values according to the user’s preferences.
Appendix A Proofs
Here we give the proofs for the theoretical work in the paper.
Proof for Proposition 4:
Proof.
[…]
Proof for Proposition 5.1:
Proof.
Let $a,b\in\mathcal{X}$ be two adjacent arguments in the ordering, such that $a\succ b$ in the preferencing ordering. According to the Definition 9 and Algorithm 1, the distance assigned to $a$ and $b$ satisfies $d_{b}=d_{a}+\delta$. Where $\delta>0$ for a regular preference ($\succ$) and $\Delta$¿0 for a much greater preference ($\succ\!\succ$). Both proposed base score extraction functions $\nu_{1}$ and $\nu_{2}$ are strictly decreasing with respect to $d$. Therefore, if $d_{a}<d_{b}$, it follows that $\nu(D,d_{a})>\nu(D,d_{b})$. This implies that the base score of the preferred argument $a$ is greater than that of $b$, i.e., $\tau(a)>\tau(b)$. Hence, setting $\delta>0$ (and similarly $\Delta>0$) ensures that both base score extraction functions satisfy Axiom 1.
Proof for Proposition 5.1:
Proof.
Let $a,b,c,d\in\mathcal{X}$ with a preference ordering where $a\succ\!\succ b\succ c\simeq d$. According to Definition 9 and Algorithm 1, the distances assigned to these arguments satisfy
$d_{b}=d_{a}+\Delta,\qquad d_{c}=d_{b}+\delta,\qquad d_{d}=d_{c},$
where $\Delta$ and $\delta$ are the increments for the relations $\succ\!\succ$ and $\succ$, respectively.
For both proposed base score extraction functions, $\nu_{1}$ and $\nu_{2}$, the base score difference between two consecutive arguments is proportional to their distance difference. Hence,
$\tau(a)-\tau(b)\propto\Delta,\qquad\tau(b)-\tau(c)\propto\delta,\qquad\tau(c)-\tau(d)=0.$
If $\Delta>\delta>0$, it follows that
$\tau(a)-\tau(b)>\tau(b)-\tau(c)>\tau(c)-\tau(d),$
which satisfies Axiom 2. Therefore, setting $\Delta>\delta>0$ guarantees that both $\nu_{1}$ and $\nu_{2}$ respect the intended magnitude differences between much greater and greater preferences.
Proof for Proposition 5.1:
Proof.
Let $\nu$ be a base score extraction function applied to a set of arguments $X$ with a preference ordering $\succeq_{p}$. By definition, $\nu_{p}(X)$ assigns base scores according to the relative ordering and distances among the arguments.
Suppose now that the preference ordering changes to a non-isomorphic ordering $\succeq_{p^{\prime}}$. By Definition 1, two preference orderings are isomorphic if and only if there exists a bijection $f:X\to X^{\prime}$ such that, for all $i,j\in X$,
$i\succeq_{p}j\;\Leftrightarrow\;f(i)\succeq_{p^{\prime}}f(j).$
If $\succeq_{p}$ and $\succeq_{p^{\prime}}$ are not isomorphic, then at least one pair of arguments $(i,j)$ exists such that their preference relations differ, i.e.,
$i\succ_{p}j\quad\text{but}\quad f(i)\not\succ_{p^{\prime}}f(j),$
or vice versa.
Because both proposed base score extraction functions $\nu_{1}$ and $\nu_{2}$ are monotonic with respect to the preference order, a change in the ordering of any pair of arguments necessarily alters their relative distances $d_{i},d_{j}$ and hence their base scores $\tau(i)$ and $\tau(j)$. Consequently,
$\nu_{p}(X)\neq\nu_{p^{\prime}}(X),$
which satisfies Axiom 3. ∎
Proof for Proposition 5.1:
Proof.
Given the definition of $\nu_{1}$, if we set $\top=1$ and $\bot=0$, then
$\nu_{1}(1,0,D,d)\;=\;0+(1-0)\,\frac{D-d}{D-1}\;=\;\frac{D-d}{D-1}.$
For the most preferred argument, we have $d=1$, hence
$\nu_{1}(1,0,D,1)\;=\;\frac{D-1}{D-1}\;=\;1.$
For the least preferred argument, we have $d=D$, hence
$\nu_{1}(1,0,D,D)\;=\;\frac{D-D}{D-1}\;=\;0.$
Because $d\in\{1,\dots,D\}$ and the function (Dd)/(D1)(D-d)/(D-1) decreases linearly from $1$ to 0 as $d$ goes from $1$ to $D$, all base scores lie in $[0,1]$ with the required endpoints. Therefore, $\nu_{1}$ with $\top=1$ and $\bot=0$ satisfies the Base Score Normalisation property. ∎
Proof for Proposition 5.1:
Proof.
Given the definition of $\nu_{1}$, if we set $\top=1$-$\bot$ with $\bot\leq 0.5$, then
$\top-\bot\;=\;(1-\bot)-\bot\;=\;1-2\bot.$
Substituting into $\nu_{1}$ yields
$\nu_{1}(1-\bot,\bot,D,d)\;=\;\bot+(1-2\bot)\,\frac{D-d}{D-1}.$
For the most preferred argument ($d=1$),
$\nu_{1}(1-\bot,\bot,D,1)\;=\;\bot+(1-2\bot)\,\frac{D-1}{D-1}=\bot+(1-2\bot)=1-\bot=\top.$
For the least preferred argument ($d=D$),
$\nu_{1}(1-\bot,\bot,D,D)\;=\;\bot+(1-2\bot)\,\frac{D-D}{D-1}=\bot.$
Moreover, the average of the endpoint values is
$\frac{\top+\bot}{2}\;=\;\frac{(1-\bot)+\bot}{2}\;=\;\frac{1}{2},$
so the endpoints are symmetric about $0.5$. Since $\bot\leq 0.5$ guarantees $\top\geq\bot$, the assignment is valid within $[0,1]$. Therefore, setting $\top=1$-$\bot$ (with $\bot\leq 0.5$) makes $\nu_{1}$ satisfy the Base Score Centralisation property. ∎
Proof for Proposition 5.1:
Proof.
Given the definition of $\nu_{2}$, if we set $\alpha=\beta=0$, then
$\nu_{2}(0,0,D,d)\;=\;\frac{D-d}{D-1}.$
For the most preferred argument, we have $d=1$, hence
$\nu_{2}(0,0,D,1)\;=\;\frac{D-1}{D-1}\;=\;1.$
For the least preferred argument, we have $d=D$, hence
$\nu_{2}(0,0,D,D)\;=\;\frac{D-D}{D-1}\;=\;0.$
Since $d\in\{1,\dots,D\}$, the value (Dd)/(D1)(D-d)/(D-1) ranges linearly from $1$ to 0 as $d$ goes from $1$ to $D$, and all values lie in $[0,1]$. Therefore, $\nu_{2}$ with $\alpha=\beta=0$ satisfies the Base Score Normalisation property. ∎
Proof for Proposition 5.1:
Proof.
Given the definition of $\nu_{2}$, for the most preferred argument ($d=1$) we obtain the upper limit
$\top\;=\;\nu_{2}(\alpha,\beta,D,1)=\frac{D-1+\alpha}{D-1+\beta},$
and for the least preferred argument ($d=D$), the lower limit is
$\bot\;=\;\nu_{2}(\alpha,\beta,D,D)=\frac{\alpha}{D-1+\beta}.$
We require $\top=1$-$\bot$. Substituting the expressions above and bringing them to a common denominator gives
$\frac{D-1+\alpha}{D-1+\beta}\;=\;1-\frac{\alpha}{D-1+\beta}\;=\;\frac{D-1+\beta-\alpha}{D-1+\beta}.$
Equating numerators yields
$D-1+\alpha\;=\;D-1+\beta-\alpha\quad\Longrightarrow\quad 2\alpha=\beta,$
i.e. $\alpha=\beta/2$. Hence, setting $\alpha=\beta/2$ ensures $\top=1$-$\bot$, so $\nu_{2}$ satisfies the Base Score Centralisation property.
Proof for Proposition 5.1:
Proof.
Let $(a,b)$ and $(c,d)$ be two pairs of adjacent arguments in a preference ordering, and suppose both pairs share the same preference relation type (i.e. both are $\succ$, both are $\succ\!\succ$, or both are $\simeq$). By Algorithm 1, the distance difference between the members of each adjacent pair is identical to the increment associated with that relation:
[…]
We show that for both proposed base score extraction functions, the base score difference between the elements of an adjacent pair depends only on this distance difference (hence, it is the same for both pairs).
For $\nu_{1}$:
$\nu_{1}(\top,\bot,D,d)=\bot+(\top-\bot)\frac{D-d}{D-1}.$
Thus, for an adjacent pair $(x,y)$ we have
$\tau(x)-\tau(y)$ $=\nu_{1}(\top,\bot,D,d_{x})-\nu_{1}(\top,\bot,D,d_{y})$
$=(\top-\bot)\frac{(D-d_{x})-(D-d_{y})}{D-1}$
$=(\top-\bot)\frac{d_{y}-d_{x}}{D-1}.$
Hence if $d_{y}-d_{x}$ is the same (e.g. $\delta$ or $\Delta$) for both pairs, then $\tau(a)-\tau(b)=\tau(c)-\tau(d)$. For the indifference relation ($\simeq$), $d_{y}-d_{x}=0$ and the difference is 0.
For $\nu_{2}$:
$\nu_{2}(\alpha,\beta,D,d)=\frac{D-d+\alpha}{D-1+\beta}.$
For an adjacent pair $(x,y)$ we obtain
$\tau(x)-\tau(y)=\frac{D-d_{x}+\alpha}{D-1+\beta}-\frac{D-d_{y}+\alpha}{D-1+\beta}=\frac{d_{y}-d_{x}}{D-1+\beta}.$
Again, this difference depends only on $d_{y}-d_{x}$; therefore, equal increments $\delta$ or $\Delta$ produce equal base score differences for the two pairs, and indifference yields zero difference.
Since fixing $\delta$ and $\Delta$ makes the distance differences between any two adjacent arguments with the same relation identical, both $\nu_{1}$ and $\nu_{2}$ assign equal base score gaps to such adjacent pairs. This is precisely the statement of Base Score Relation Regularity. ∎
Proof for Proposition 5.1:
Proof.
Let $\succeq_{p}$ be a preference ordering over $\mathcal{X}$ and let $S$ be a set of new arguments such that $\mathcal{X}^{\prime}=\mathcal{X}\cup S$ and for every $s\in S$ there exists $a\in\mathcal{X}$ with $a\simeq s$ under the extended ordering $\succeq_{p^{\prime}}$. By construction, adding $S$ introduces only indifference relations ($\simeq$) between new arguments and some existing arguments in $\mathcal{X}$. Observe two points about Algorithm 1 and the distance assignment (Definition 9):
  1. (1)
    The total distance span $D=1+n\cdot\delta+m\cdot\Delta$ depends only on the counts $n$ and $m$ of $\succ$ and $\succ\!\succ$ relations in the ordering. Since every $s\in S$ is indifferent to some $a\in\mathcal{X}$, the numbers $n$ and $m$ for relations between distinct preference tiers remain unchanged; hence $D$ is unchanged.
  2. (2)
    When traversing the ordering, Algorithm 1 increases the running distance $d$ only when the previous relation was $\succ$ (by $\delta$) or $\succ\!\succ$ (by $\Delta$), and leaves $d$ unchanged when the previous relation is $\simeq$. Therefore, inserting arguments that are indifferent to existing ones does not change the distance values assigned to the original arguments in $\mathcal{X}$ (they either share the same distance as before or receive the same $d$ as their indifferent partner).
Because both $\nu_{1}(\top,\bot,D,d)$ and $\nu_{2}(\alpha,\beta,D,d)$ depend only on $D$ and the per-argument distances $d$, and neither $D$ nor the distances of original arguments change under the described insertion of indifferent arguments, it follows that the base scores assigned to arguments in $X$ remain identical:
$\nu_{p}(\mathcal{X})\;=\;\nu_{p^{\prime}}(\mathcal{X}^{\prime}\setminus S).$
Hence, both $\nu_{1}$ and $\nu_{2}$ satisfy Base Score Preference Stability. ∎
Proof for Proposition 5.1:
Proof.
Let $\nu$ be a base score extraction function satisfying Axiom 1, and let $\sigma$ be a gradual semantics that satisfies monotonicity and balance.
Take two arguments $a,b\in X$ such that $a\succ b$ and suppose that $a$ and $b$ have equivalent sets of supporters and attackers (i.e. the multiset of incoming supports and attacks and their sources are identical up to renaming between $a$ and $b$). By Axiom 1, we have $\tau(a)>\tau(b)$.
Now consider the evaluation of $\sigma$ on the QBAF where all base scores except those of $a$ and $b$ are held fixed. Monotonicity of $\sigma$ implies that, when all other inputs are unchanged, increasing an argument’s base score (while keeping others fixed) cannot decrease its final strength:
$\tau(a)>\tau(b)\quad\Longrightarrow\quad\sigma(a)\geq\sigma(b),$
and more precisely, a strict increase in the base score (with identical influences) yields a strict increase in the final strength under standard monotonicity assumptions used in the literature.
Balance ensures that the semantics treats arguments that have the same structural influences (same attackers/supporters with the same base scores) symmetrically: if two arguments have identical incoming influences and identical incoming relations, then differences in their final strengths are entirely attributable to differences in their own base scores.
Combining these two properties: since $a$ and $b$ receive the same aggregated influence from the rest of the framework (by hypothesis) and $\tau(a)>\tau(b)$ (by Axiom 1), the monotonicity and balance of $\sigma$ imply
$\sigma(a)>\sigma(b).$
Thus, when $a\succ b$ and $a,b$ have equivalent supporters and attackers, the semantics $\sigma$ ranks $a$ strictly above $b$. ∎